OpenStudy (jakeybabs):

19-x-x=-13 whats the answer for x?

OpenStudy (anonymous):

Simplifying: 19 + -1x + -1x = -13 Combine like terms: -1x + -1x = -2x 19 + -2x = -13 Solving: 19 + -2x = -13 Move all terms containing x to the left, all other terms to the right. Add '-19' to each side of the equation. 19 + -19 + -2x = -13 + -19 Combine like terms: 19 + -19 = 0 0 + -2x = -13 + -19 -2x = -13 + -19 Combine like terms: -13 + -19 = -32 -2x = -32 Divide each side by '-2'. x = 16 Simplifying: x = 16

OpenStudy (jakeybabs):

thank you ver much :)

OpenStudy (anonymous):

You're welcome :)